OpenStudy (anonymous):

What's the formula for finding the area and perimeter of a right triangle???

OpenStudy (anonymous):



OpenStudy (anonymous):

area=(1/2)*a*b and perimeter =a+b+H

OpenStudy (anonymous):

what if C is blank??

OpenStudy (anonymous):

i didnt understand the ques

OpenStudy (anonymous):

what do you mean :D

OpenStudy (anonymous):

what if there is no number for c?